# Towards Autonomous Driving with Small-Scale Cars: A Survey of Recent Development

### v1

@@ Abstract
While engaging with the unfolding revolution in autonomous driving, a challenge presents itself, how can we effectively raise awareness within society about this transformative trend? While full-scale autonomous driving vehicles often come with a hefty price tag, the emergence of small-scale car platforms offers a compelling alternative. These platforms not only serve as valuable educational tools for the broader public and young generations but also function as robust research platforms, contributing significantly to the ongoing advancements in autonomous driving technology. This survey outlines various small-scale car platforms, categorizing them and detailing the research advancements accomplished through their usage. The conclusion provides proposals for promising future directions in the field.
@@
Small-scale car, autonomous vehicles, robotics
@@ I Introduction
Over the past few decades, extensive research on autonomous driving (AD) has been conducted by both academic communities and industry stakeholders. However, the fruition of fully functional applications for AD systems is anticipated to materialize in the coming decades. The question arises: Can we proactively prepare our society for the oncoming fully autonomous driving? Despite assertions from researchers that autonomous vehicles (AVs) will mitigate human error and enhance safety compared to human drivers, public apprehensions persist regarding the ethical and safety dimensions of AVs [1, 2]. Encouragingly, individuals with prior experience with AVs and younger generations exhibit a more optimistic stance toward these technologies [3]. Hence, the affirmative response to this question is clear. Given that those who will design autonomous systems are currently in high school, and those who will coexist with AVs are presently in elementary schools, the optimal preparation involves providing opportunities for the public to have touch with AD systems now. To this end, a small-scale car platform emerges as an ideal choice, serving not only educational purposes but also facilitating researchers in testing their autonomous systems on a tangible platform.
Undoubtedly, research in AD research for full-scale cars holds promise. However, to provide the public and research communities with a more cost-effective and accessible avenue to engage with AD, the replication of small-scale car platforms stands out. In educational settings, small-scale cars prove to be an excellent choice for schools, especially with STEM focus, offering students their initial exposure and hands-on experience with AD while simultaneously fostering increased public awareness. On the research front, the availability of such research platforms lowers entry barriers, inviting a broader spectrum of researchers into AD exploration. This, in turn, propels advancements in AD and establishes a virtuous circle of continuous improvement and innovation before their integration into full-scale autonomous vehicles. As evidenced in Fig. 1(A), the quantity of research papers focused on small-scale car platforms has experienced a substantial surge in the past two decades. Following a relatively gradual incline before 2016, this field has undergone a notable expansion, particularly catalyzed by the introduction of a series of well-known platforms. Hence, to advance ongoing research initiatives and enhance public awareness, we attempt to comprehensively survey the existing small-scale platforms for educational and research purposes. We also aim to benchmark the AD tasks accomplished by these platforms. To the best of our knowledge, this survey represents the first of its kind to undertake such an extensive analysis.
@@
The rest of this paper is organized as follows: In the second section, we examine commonly employed small-scale car platforms, accompanied by insights into their simulators. The third section provides a comprehensive overview of the AD tasks accomplished within the research community. Following that, the fourth section inspects the sensor setups employed in these platforms. The subsequent section outlines promising directions for future developments, and we draw our conclusions in the sixth section.
@@
Figure 1: Illustration of the development and current states of small-scale car platforms, each depicted platform image sourced from its respective paper or website. (A) Based on published studies each year on Google Scholar with the search terms: robot car, the research on small-scale cars has seen substantial growth in the number of papers over the years. In the early 2000s, projects like s-bot [4], e-puck [5], and TurtleBot [6] emerged. Starting around 2016, with the introduction of Duckietown [7], BARC [8], and Autorally [9], there was a significant increase in research papers. This trend continued with the development of projects like DeepRacer [10], Donkeycar [11], and F1TENTH [12]. More computationally advanced small-scale cars have been introduced in recent years, such as ART/ATK [13] and XTENTH-CAR [14]. (B) Examples of small-scale car platforms, categorized into educational platforms and research platforms, including multiple vehicle setups such as ORAC [15] and UDSSC [16].
@@ II Platforms
For full-scale cars, the standardized dynamics and sizes are mandated by road regulations. In contrast, smaller-scale cars, designed without regulatory constraints, are crafted for diverse research purposes by various groups of researchers and enthusiasts. Recognizing these disparities, this survey establishes a clear definition for AD robot cars, setting them apart from other autonomous robots. In this context, an AD small-scale car platform refers to a miniature vehicle equipped with technologies enabling autonomous operation. This includes sensors, processing units, control actuators, and often a set of predefined tasks or challenges for testing and development. In terms of size, namely the scale factors, there is considerable variability among different platforms. Platforms intended to carry a lot of sensors and for outdoor operation can reach almost 1 meter in length [9]. On the other end of the spectrum, platforms developed for investigating swarm dynamics are notably compact, with some as small as 3.3x3.3cm [17]. The most common form factor is the 1/10th scale, representing a model approximately one-tenth the size of a full-scale car. This scale is frequently employed in hobby Remote-Control (RC) vehicles, forming the foundation for numerous platforms [11, 12, 18, 8, 19, 20, 21, 22, 23, 24, 14]. Regarding the dynamic system, full-size cars commonly use Ackermann steering mechanics, rotating the front axle to facilitate left or right turns [25]. However, due to the mechanical complexity of this system and its limitations in certain situations, small-scale car platforms initially favored the use of a differential steering system. In this system, the left and right sets of wheels are driven independently, enabling turns by driving one set of wheels faster than the other. Nevertheless, as illustrated in Fig. 1, recent advancements in this field have seen the increased adoption of the more realistic Ackermann steering system in small-scale car platforms.
@@ II-A Hardware platforms
Before we introduce the different platforms, we categorize small-scale car platforms into distinct groups based on their target users and complexity. First are the educational platforms often powered by MicroBit or Raspberry Pi, predominantly accessible in the commercial market, including Makeblock mBot [26], Edison [27], AlphAI Robot [28], TinkerGen MARK [29], and Robolink Zumi [30], typically using platform-specified block and visual programming languages. These platforms are carefully crafted to engage students ranging from early childhood to elementary school and even graduate programs in the interactive exploration of autonomous cars. They offer users comprehensive tools, enabling them to effortlessly construct and program robot cars capable of executing diverse tasks. With these platforms, students can readily accomplish basic tasks like simple lane following and obstacle avoidance.
@@
The subsequent tier has a more sophisticated system, affording users enhanced opportunities to innovate and develop new functionalities within the platform. The typical platforms are e-puck [5], Pheeno [31], Thymio [32], MarXbot [33], Turtlebot [6], and Duckietown [7]. The e-puck, Pheeno, Thymio, and MarXbot are commonly employed in swarm robot research due to their compact sizes. The Duckietown ecosystem, originating in 2016 at MIT, serves as an educational tool for instructing students in the realms of autonomous and connected driving. Comprising cost-effective vehicles known as Duckiebots, each unit is outfitted with either a Raspberry Pi or a Jetson Nano as its computing unit. Sensor-wise, the Duckiebots employ an RGB Raspberry Pi camera, an inertial measurement unit (IMU), wheel encoders, and a front-facing distance sensor. The platform also facilitates the incorporation of a central localization system through ”watchtowers.” Propelled by two DC motors on each of the two driven wheels, the cars adhere to a differential drive pattern. Duckietown is commercially available and has proven effective in various applications, with lane-keeping using the RGB camera being the most common [34, 35]. Additionally, it has been employed for tasks such as obstacle and traffic sign detection [36, 37]. Another commercially available platform is the AWS DeepRacer [10] by Amazon Web Services. This platform also relies on an RGB camera for sensory input, supplemented by an upgrade kit featuring a second camera and a LiDAR sensor. The execution of learned strategies is managed by a small compute board equipped with an Intel Atom CPU. With Ackermann steering implementation, DeepRacer is also primarily used for tasks like lane-keeping [38, 39] and obstacle avoidance [40, 41].
For users seeking open-source platforms and aiming to construct a platform from the ground up, two common options are Donkeycar [11] and F1TENTH [12]. Donkeycar, a collaborative effort within a community, provides comprehensive instructions for constructing the platform. Due to its open-source nature, individual cars can vary significantly, accommodating diverse sensor configurations. The foundational platform centers around a 1/10th scale RC car, featuring Ackermann steering, and is customized with a Raspberry Pi or Jetson Nano as the compute unit, along with a Raspberry Pi camera serving as its primary sensor. Users have the flexibility to incorporate additional sensors such as IMUs and encoders, and there are options for integrating LiDAR into the system. Given its open-source framework, users can extend the platform’s functionality according to their needs. Originally designed for autonomous racing, Donkeycar is also commonly applied for tasks like lane-keeping [42, 43] and obstacle avoidance [44]. Likewise, F1TENTH is built upon a 1/10th scale RC car platform featuring Ackermann steering. Its compute unit is a Nvidia Jetson, complemented by a 2D LiDAR serving as the primary sensor. Additional features include an IMU and odometry information provided by the VESC motor controller [^1] https://vesc-project.com/, with an option to integrate an extra RGB camera. With its high-speed capable base platform, F1TENTH is primarily utilized for autonomous racing tasks [45, 46], while, it also finds application in fundamental tasks such as lane-keeping [47] and obstacle avoidance [48, 49].
For outdoor applications, the Autorally [9] platform takes center stage, incorporating an off-the-shelf RC car platform. In comparison to the Donkeycar and F1TENTH platforms, Autorally is larger, utilizing a 1/5th scale model car. Leveraging this increased size, it incorporates a more powerful compute unit consisting of a consumer computer mainboard with an Intel i7 CPU. The primary sensors include two industrial synchronized RGB cameras for obtaining stereo images, an IMU, magnetic encoders for odometry measurement, and a GPS receiver. Due to its substantial size and high-performance components, Autorally is mainly employed for racing tasks [50, 51, 52]. The platform offers a variety of sensor combinations for racing approaches, with many leveraging the combination of GPS and IMU data [53, 54], as well as those that integrate camera and IMU data [55, 56].

### v2

# Autonomous Driving Small-Scale Cars: A Survey of Recent Development

@@ Abstract
While engaging with the unfolding revolution in autonomous driving, a challenge presents itself, how can we effectively raise awareness within society about this transformative trend? While full-scale autonomous driving vehicles often come with a hefty price tag, the emergence of scaled-down, small-scale car platforms offers a compelling alternative. These miniature vehicles are designed to perform predefined tasks and challenges, equipped with onboard sensors, processing units, and control actuators. These platforms not only serve as valuable educational tools for the broader public and young generations but also function as robust research platforms, contributing significantly to the ongoing advancements in autonomous driving technology. This survey outlines various small-scale car platforms, categorizing them and detailing the research advancements accomplished through their usage. The conclusion provides proposals for promising future directions in the field.
@@
Small-scale car, autonomous driving, robotics
@@ I Introduction
Over the past few decades, extensive research on autonomous driving (AD) has been conducted by both academic communities and industry stakeholders. However, the fruition of fully functional Level 5 autonomous driving systems, which entails the availability of completely autonomous vehicles in the mass market, is anticipated to materialize in the coming decades [1, 2]. The question arises: Can we proactively prepare our society for the oncoming fully autonomous driving? Despite assertions from researchers that autonomous vehicles (AVs) will mitigate human error and enhance safety compared to human drivers, public apprehensions persist regarding the ethical and safety dimensions of AVs [3, 4]. Encouragingly, individuals with prior experience with AVs and younger generations exhibit a more optimistic stance toward these technologies [5]. Hence, the affirmative response to this question is clear. Given that those who will further develop the AD systems are currently in high school, and those who will coexist with AVs are presently in elementary schools, the optimal preparation involves providing opportunities for the public to engage with AD systems now. To this end, a small-scale car platform emerges as an ideal choice, serving educational purposes and facilitating researchers in testing their autonomous systems on a tangible platform.
Undoubtedly, research in AD research for full-scale cars holds promise. Despite the supportive environments for developing and testing AVs in various countries and regions, strict regulations persist due to concerns about safety, security, and public trust [6, 7, 8]. These regulations often limit the ability to drive or test AVs on public roads, which consequently restricts the involvement of smaller research institutions or individual researchers in AD system development. As a solution, the use of small-scale car platforms offers a more cost-effective and accessible alternative for the public and research communities to engage with AD technologies. More specifically, small-scale car platforms can serve two main purposes: educational tools for students and research tools for AD researchers. In educational settings, small-scale cars prove to be an excellent choice for schools, especially with science, technology, engineering, and mathematics (STEM) focus, offering students their initial exposure and hands-on experience with AD while simultaneously fostering increased public awareness. On the research front, the availability of such research platforms lowers entry barriers, inviting a broader spectrum of researchers into AD exploration. As evidenced in Fig. 1(A), based on published studies each year on Google Scholar with the search terms: ”small-scale car” or ”robot car” from 2000 onward, the quantity of research papers focused on small-scale car platforms has experienced a substantial surge in the past two decades. Following a relatively gradual incline before 2016, this field has undergone a notable expansion, particularly catalyzed by the introduction of a series of well-known platforms. It is worth mentioning that AD research for small-scale cars is the ultimate goal of these small-scale car platforms, not the transformation of the techniques into real-size vehicles. Various major small-scale car platforms claim that these platforms are designed to be accessible and inexpensive, aimed at fostering educational and research activities [9, 10, 11, 12, 13]. For example, as stated by Duckietown platform [9], their mission is Learning robotics and AI made tangible, accessible, and fun! They want to provide a tangible, accessible, and inclusive tool for a broader range of society to engage with robots and AV. Similarly, another major platform DeepRacer [10], provided by Amazon Web Services, also tries to offer developers of all skill levels for hands-on experience with fully autonomous 1/18th scale race cars with machine learning (ML). Another platform, F1TENTH’s mission is to foster interest, excitement, and critical thinking about the increasingly ubiquitous field of autonomous systems [12]. In comparison to full-scale AVs, small-scale car platforms employ more lightweight computation units and sensors for economic reasons. These platforms primarily use mature AV techniques from full-scale cars, albeit under simplified conditions due to current hardware limitations. Nonetheless, they play a crucial role in engaging undergraduate students and early career researchers in AV technologies, inspiring them, and contributing to the growth of the AV research community. This, in turn, indirectly advances real-world AV research.
I-A Contributions
Despite the importance and widespread use of small-scale car platforms, no research currently provides a comprehensive overview of these platforms, including their hardware configurations, software frameworks, and benchmarking of driving tasks. In [14], the existing literature on small-scale cars is summarized; however, it lacks comprehensiveness and does not include a comparative analysis. More recently, [15] discussed existing platforms, but their focus was limited to the platforms themselves, without addressing the techniques used to advance small-scale car AD research. Identifying this gap, we aim to advance ongoing research initiatives and enhance public awareness by conducting a comprehensive survey of existing small-scale platforms for educational and research purposes. Additionally, we seek to benchmark the AD tasks accomplished by these platforms. To the best of our knowledge, this survey represents the first of its kind to provide such an extensive analysis. The contributions of this work are as follows:
  • Comprehensive Review of Platforms: We thoroughly review the most widely available small-scale car platforms, comparing their hardware and software configurations to provide readers with an in-depth understanding and a guide for selecting or building their own platforms.
  • Benchmarking Autonomous Driving Tasks: By analyzing over 250 research papers, we categorize the driving tasks for small-scale cars into two primary pipelines—modular pipelines and end-to-end pipelines. We discuss in detail the techniques employed in each module within these pipelines.
  • Proposed Future Directions: After examining the existing techniques, we propose potential improvements for both the software and hardware aspects. On the software side, we suggest techniques to enhance driving behaviors, while on the hardware side, we recommend platform improvements to facilitate more advanced research.
@@ I-A Contributions
The rest of this paper is organized as follows: In Section II, we examine commonly employed small-scale car platforms, accompanied by insights into their simulators. Following that in Section III, we inspect the sensor setups employed in these platforms. Section IV provides a comprehensive overview of the AD tasks accomplished within the research community. Section V outlines promising directions for future developments, and we draw our conclusions in Section VI.
@@
Figure 1: Illustration of the development and current states of small-scale car platforms, each depicted platform image sourced from its respective paper or website. (A) Based on published studies each year on Google Scholar with the search terms: ”small-scale car” or ”robot car”, the research on small-scale cars has seen substantial growth in the number of papers over the years. In the early 2000s, projects like s-bot [16], e-puck [17], and TurtleBot emerged. Starting around 2016, with the introduction of Duckietown [9], BARC [13], and Autorally [11], there was a significant increase in research papers. This trend continued with the development of projects like DeepRacer [10], Donkeycar, and F1TENTH [12]. More computationally advanced small-scale cars have been introduced in recent years, such as ART/ATK [18] and XTENTH-CAR [19]. (B) Examples of small-scale car platforms, categorized into educational platforms and research platforms, including multiple vehicle setups such as ORAC [20] and UDSSC [21].
@@ II Platforms
For full-scale cars, the standardized dynamics and sizes are mandated by road regulations. In contrast, smaller-scale cars, designed without regulatory constraints, are crafted for diverse research purposes by various groups of researchers and enthusiasts. Recognizing these disparities, this survey establishes a clear definition for AD robot cars, setting them apart from other autonomous robots. In this context, an AD small-scale car platform refers to a miniature vehicle equipped with technologies enabling autonomous operation. This includes sensors, processing units, control actuators, and often a set of predefined tasks or challenges for testing and development. In terms of size, namely the scale factors, there is considerable variability among different platforms. Platforms intended to carry a lot of sensors and for outdoor operation can reach almost 1 meter in length [11]. On the other end of the spectrum, platforms developed for investigating swarm dynamics are notably compact, with some as small as 3.3x3.3cm [22]. The most common form factor is the 1/10th scale, representing a model approximately one-tenth the size of a full-scale car. This scale is frequently employed in hobby Remote-Control (RC) vehicles, forming the foundation for numerous platforms [23, 12, 24, 13, 25, 26, 27, 28, 19]. Regarding the dynamic system, full-scale cars commonly use Ackermann steering mechanics, rotating the front axle to facilitate left or right turns [29]. However, due to the mechanical complexity of this system and its limitations in certain situations, small-scale car platforms initially favored the use of a differential steering system. In this system, the left and right sets of wheels are driven independently, enabling turns by driving one set of wheels faster than the other. Nevertheless, as illustrated in Fig. 1, recent advancements in this field have seen the increased adoption of the more realistic Ackermann steering system in small-scale car platforms.
@@ II-A Hardware Platforms
Before we introduce the different platforms, we categorize small-scale car platforms into distinct groups based on their target users and complexity.
II-A1 Educational Platforms
First are the educational platforms predominantly accessible in the commercial market, as shown in Table I, including Makeblock mBot, Edison, AlphAI Robot, TinkerGen MARK, and Robolink Zumi. These platforms are carefully crafted to engage students ranging from early childhood to elementary school and even graduate programs in the interactive exploration of autonomous cars. They offer users comprehensive tools, enabling them to construct effortlessly and program robot cars capable of executing diverse tasks. These platforms are equipped with basic sensors such as line sensors, distance sensors, and occasionally cameras. With limited processing power, most rely on microcontrollers, they are suitable only for simple tasks like lane keeping or car following. However, their support for simple frameworks and visual programming languages makes them ideal for undergraduate education. Many also come with pre-made courses designed to teach students the fundamentals of their operation. Their low cost and commercial availability further enhance their accessibility for educational purposes. Despite these advantages, these platforms are not well-suited for researchers exploring autonomous or connected driving. Their closed-source hardware and software limit extensibility, and most lack networking capabilities for external control or integration into multi-vehicle scenarios. Additionally, their differential-drive steering design does not accurately replicate real car dynamics, making them inadequate for serious investigations into realistic vehicular behavior.
Platforms Size Sensors Programming Runtime Price (USD)
Thymio [30] 110x110mm IR sensor, Accelerometer, Microphone, Thermistor VPL 2h 270
Makeblock mBot Ultrasonic Sensor, IR sensor, Line tracking sensor mBlock 1h 190
Edison 80x80x40mm IR sensor, Line tracking sensor EdBlock EdPy 60
AlphAI Robot Camera, Ultrasonic Sensor, Line tracking sensor Python 270
Ozobot Evo 32mm ⌀ IR sensor, Speaker OzoBlockly 1h 160
Finch Robot Encoder, Ultrasonic sensor, Line tracking sensor, IR sensor, Speaker MicroBit 7h 170
TinkerGen MARK 200x185x92mm Camera, Ultrasonic Sensor MicroPython 220
Robolink Zumi 95x67x70mm Camera, IR sensor Python 190
TABLE I: An overview of small-scale car platforms for educational purposes.
II-A2 Commercial Research Platforms
@@
The subsequent tier has a more sophisticated system, affording users enhanced opportunities to innovate and develop new functionalities within the platform. These platforms also mostly offer support for more advanced frameworks like Robot Operating System (ROS) and standard programming languages such as Python to incorporate them in bigger research projects. Their inclusion of networking capabilities also aids in incorporating them in diverse environments. The typical platforms include e-puck [17], Pheeno [31], Thymio [30], MarXbot [32], Turtlebot, and Duckietown [9]. The e-puck, Pheeno, Thymio, and MarXbot are frequently used in swarm robotics research due to their compact size and reliance on basic sensors. The Duckietown ecosystem originated in 2016 at MIT as an educational tool for instructing students in the realms of autonomous and connected driving but developed into a research platform for all levels of researchers. Comprising cost-effective vehicles known as Duckiebots, each unit is outfitted with either a Raspberry Pi or a Jetson Nano as its computing unit. Sensor-wise, the Duckiebots employ an RGB Raspberry Pi camera, an inertial measurement unit (IMU), wheel encoders, and a front-facing distance sensor. The platform also facilitates the incorporation of a central localization system through ”watchtowers.” Propelled by two DC motors on each of the two driven wheels, the cars adhere to a differential drive pattern. Duckietown is commercially available and widely used for tasks such as lane keeping using the RGB camera [33, 34], as well as obstacle and traffic sign detection [35, 36]. Its support for diverse frameworks and sensors, along with its open-source nature, makes it applicable to various target groups and tasks. Comprehensive documentation and a large community further enhance its adaptability and integration into different environments. The differential-drive steering setup in Duckietown, similar to educational platforms, is a significant limitation. Another commercially available platform is the AWS DeepRacer [10] by Amazon Web Services. This platform also relies on an RGB camera for sensory input, supplemented by an upgrade kit featuring a second camera and a LiDAR sensor. The execution of learned strategies is managed by a small compute board equipped with an Intel Atom CPU. With Ackermann steering implementation, DeepRacer is also primarily used for tasks like lane keeping [37, 38] and obstacle avoidance [39, 40]. While the platform offers open-source software to interface with its hardware, its heavy reliance on Amazon Web Services limits extensibility beyond the intended use of the manufacturer. However, AWS provides extensive documentation, including detailed courses on setup and operation within its software ecosystem.
II-A3 Open Source Platforms
For users seeking open-source platforms and aiming to build a system from the ground up, two popular options are Donkeycar[23] and F1TENTH [12]. Donkeycar is a community-driven platform that provides detailed instructions for constructing a customizable vehicle. Its open-source nature allows significant flexibility in sensor configurations and system design. The base platform is built on a 1/10th-scale RC car with Ackermann steering, equipped with a Raspberry Pi or Jetson Nano as the compute unit and a Raspberry Pi camera as the primary sensor. Users can enhance the platform by adding sensors such as IMUs, encoders, and even LiDAR. Given its open-source framework, users can extend the platform’s functionality according to their needs. Originally designed for autonomous racing, Donkeycar is commonly used for tasks like lane keeping [41, 42] and obstacle avoidance [43]. However, its versatility extends to a wide range of applications due to its open-source framework. The community behind the project also offers comprehensive documentation on building and using the platform. Its Python-based custom framework supports integration with other systems, such as ROS, further expanding its functionality. While primarily designed for single-vehicle use, the platform could be extended to enable coordination among multiple vehicles. Despite its advantages, the complexity and the need for manual assembly and configuration may deter researchers seeking out-of-the-box solutions. F1TENTH, like Donkeycar, is based on a 1/10th-scale RC car with Ackermann steering. It features a Nvidia Jetson as its compute unit and a 2D LiDAR as the primary sensor. Additional hardware includes an IMU, odometry data provided by the VESC motor controller, and an optional RGB camera. With its high-speed capable base platform, F1TENTH is primarily utilized for autonomous racing tasks [44, 45], though it is also applied in fundamental tasks like lane keeping [46] and obstacle avoidance [47, 48]. Thanks to its relatively powerful compute unit, F1TENTH can support advanced applications, including ML-based decision-making. The big advantage of these two platforms lies in their open-source nature and comprehensive online documentation, enabling extensive customization, such as incorporating additional sensors or connecting to simulators. However, this flexibility comes with the challenge of increased complexity, as users must build and configure the platforms from scratch, an approach that may not be suitable for all researchers.
II-A4 Outdoor Platforms
For outdoor applications, the Autorally [11] platform takes center stage, built on an off-the-shelf RC car platform. Unlike smaller platforms such as Donkeycar and F1TENTH, AutoRally uses a larger 1/5th-scale model car, allowing it to accommodate more powerful hardware. Its compute unit is a consumer-grade computer mainboard equipped with an Intel i7 CPU, enabling advanced processing capabilities. The platform features a sophisticated sensor suite, including two industrial synchronized RGB cameras for stereo imaging, an IMU, magnetic encoders for odometry, and a GPS receiver. Due to its substantial size and high-performance components, Autorally is mainly employed for racing tasks [49, 50, 51]. These tasks often utilize various sensor fusion strategies, such as combining GPS and IMU data [52, 53] or integrating camera and IMU data [54, 55]. While the diverse set of sensors and substantial computing power allow the platform to be used for different tasks, its large size and high-speed capabilities necessitate an outdoor environment, making it unsuitable for smaller lab setups. The project provides instructions for constructing the vehicle, setting up the software, and integrating it into the ROS framework. However, the lack of comprehensive documentation may present challenges for users attempting to fully leverage its potential.
II-A5 Multi Vehicle Platforms